OpenStudy (anonymous):

Part 1: Change 3x + 2y = 10 to slope-intercept form. You must show all work. (2 points) Part 2: State the slope and y-intercept (2 points)

OpenStudy (anonymous):

Part 1: Y= mx +b m= slope and b is the y intercept) 3x-2y=10 make y the subject i.e. subtract 3x from both sides of the equation -2y = 10 - 3x => y = 3x/2-10/2 = 3x/2 -5 slope = 3/2 and the y intercept =-5

OpenStudy (anonymous):

Part 1: Y = 3x/2-10/2 = 3x/2 -5 Part 2: slope = 3/2 and the y intercept =-5

Directrix (directrix):

3x + 2y = 10 2y = 10 - 3x 2y = -3x + 10 y = (-3/2) x + 5 --> slope = (-3/2) and y-intercept = 5

Directrix (directrix):

@chemwile ---> one of us has a sign error on this problem. Please check my work. Thanks.